OpenStudy (anonymous):

Medal & Fan Which exponential function goes through the points (1, 8) and (4, 512)?

OpenStudy (anonymous):

f(x) = 4(2)x f(x) = 8(2)-x f(x) = 4(4)-x f(x) = 2(4)x

myininaya (myininaya):

Let's use the form \[y=a \cdot 2^{bx}\] to find out. Enter in the points you have to form two equations.

myininaya (myininaya):

Then solve the system of equations.

OpenStudy (anonymous):

whats A?

myininaya (myininaya):

idk yet that is what we have to find

myininaya (myininaya):

also have to find b

myininaya (myininaya):

to find a and b you need to form those two equations I talked about

myininaya (myininaya):

can i see your two equations?

OpenStudy (anonymous):

Im not understanding , if we dont know A or B , how can i substitute anything into that equation?

myininaya (myininaya):

you have the ordered pairs

myininaya (myininaya):

or if you don't like to plug in ordered pairs and solve a system of equations you could just take the points in enter it into your choices to see which choice has both ordered pairs on the curve

myininaya (myininaya):

\[512=a \cdot 2^{b \cdot 4 } \\ 8=a \cdot 2^{b}\] I just plugged in the ordered pairs (1,8) and (4,512) i have two equations in terms of two variables i would try solving the system by substitution you know like solve that second equation by 2^b and plug it into the first equation

myininaya (myininaya):

@briannamooney you can continue that route or since you have choices you can see which equation works for you by plugging in your ordered pairs